OpenStudy (anonymous):

Water is flowing into a spherical tank with 6 foot radius at the constant rate of 30pi cu ft per hour. When the water is h feet deep, the volume of water in the tank is given by V=(pih^2)/3*(18-h). What is the rate at which the depth of the water in the tank is increasing at the moment when the water is 2 feet deep? Answer: 1.5 ft per hr Please explain how to come to this answer.

OpenStudy (anonymous):

\[given~\frac{ dV }{dt} =30\pi ft^3/hr \] \[V=\frac{ \pi h^2 }{3 }*\left( 18-h \right) =\frac{ \pi }{3 }\left( 18 h^2-h^3 \right)\] \[\frac{ dV }{ dt }= \frac{ \pi }{3 }\left( 36h-3h^2 \right)\frac{ dh }{dt }=\pi \left( 12h-h^2 \right)\frac{ dh }{ dt }\] h=2 ft \[substitute~the~values~of~h~and~\frac{ dV }{dt }\]

OpenStudy (anonymous):

Thanks! :)

OpenStudy (anonymous):

yw